express all numbers as percentages of purchase price, and assume no residual value. For example, in year 1, depreciation expense is 20 percent and the ending book value is 80 percent. Complete for years 2 through 5.
Answer #1

Let the purchase cost be 100% .Under straight line method ,Depreciation expense = 1/ useful life

                                                                                                                            = 1/5

                                                                                                                            = .20 or 20%

YEAR DEPRECIATION EXPENSE BOOK VALUE
0 100%
1 20% 100-20 =80%
2 20% 80-20=60%
3 20% 60-20 =40%
4 20% 40-20 =20%
5 20% 20-20 =0